# This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible

### v3

@@ Abstract
This research addresses the critical lack of comprehensive studies on feature scaling by systematically evaluating 12 scaling techniques - including several less common transformations - across 14 different Machine Learning algorithms and 16 datasets for classification and regression tasks. We meticulously analyzed impacts on predictive performance (using metrics such as accuracy, MAE, MSE, and R²) and computational costs (training time, inference time, and memory usage). Key findings reveal that while ensemble methods (such as Random Forest and gradient boosting models like XGBoost, CatBoost and LightGBM) demonstrate robust performance largely independent of scaling, other widely used models such as Logistic Regression, SVMs, TabNet, and MLPs show significant performance variations highly dependent on the chosen scaler. This extensive empirical analysis, with all source code, experimental results, and model parameters made publicly available to ensure complete transparency and reproducibility, offers model-specific crucial guidance to practitioners on the need for an optimal selection of feature scaling techniques.
@@ I Introduction
Machine Learning progress has been notorious in several domains of knowledge engineering, notably driven by the rise of big data [1, 2], its applications in healthcare [3, 4, 5, 6], forecasting [7, 8, 9, 10], precision agriculture [11], wireless sensor networks [12], language tasks [13, 14, 15] and many other domains [16, 17, 18, 19, 20, 21]. All different applications compose the field of Machine Learning [14], which has become a major subarea of computer science and statistics due to its crucial role in the modern world [22]. Although these methods hold immense potential for the advancement of predictive modeling, their improper application has introduced significant obstacles [23, 24, 25].
@@
Consequently, applying a scaling method without a careful evaluation of its suitability for the specific problem and model may be inadvisable and could negatively impact results. This practice risks undermining the validity of the claims regarding model performance and may lead to a feedback loop of overconfidence in the results, known as overfitting [30].
@@
The primary objective of this study is to evaluate the impact of different data scaling methods on the training process and performance metrics of various Machine Learning algorithms across multiple datasets. We employ 14 widely used Machine Learning models for tabular data, including Linear Regression, Logistic Regression, Support Vector Machines (SVM), K-Nearest Neighbors, Multilayer Perceptron, Random Forest, TabNet, Naive Bayes, Classification and Regression Trees (CART), Gradient Boosting Trees, AdaBoost, LightGBM, CatBoost, and XGBoost. These models were evaluated using 12 different data scaling techniques, in addition to a baseline without scaling, across 16 datasets covering both classification and regression tasks. The selected models represent the state of the art in tabular data analysis, offering a favorable balance between predictive performance and computational efficiency, often outperforming deep learning techniques in this context [37, 38, 39, 40].
In Section II we cover some related work in similar studies while Section III explains more about each algorithm, each feature scaling technique, the study diagram, evaluation metrics, and how the models were trained. In Section IV we represent the final results of this study and some discussion. The limitations of our study are discussed in Section V. Lastly, in Section VI we give a final conclusion of the current experiments and future works.
@@ II Related Work
Despite its fundamental role in Machine Learning pipelines, the impact of feature scaling remains an underexplored area in the literature. Most existing studies examine only a limited number of algorithms [41] or datasets, and often provide minimal analysis of the specific effects of different scaling techniques [42] on each particular algorithm’s performance. Studies that comprehensively evaluate various scaling methods across a broad range of models and datasets, such as the approach taken in this work, are scarce. In many Machine Learning cases, preprocessing is briefly mentioned, with scaling treated as a routine step rather than a variable worthy of in-depth investigation.
@@
For some Machine Learning models, feature scaling is extremely necessary, such as K-Nearest Neighbors [43], Neural Networks [44, 45] [46], and SVM [47] [48]. In object detection, data scaling has a crucial impact [49].
In [50], the authors compared six normalization methods in an SVM classifier to improve intrusion data, and the Min-Max Normalization showed the best performance. In [51], the authors evaluated eleven Machine Learning algorithms, including Logistic Regression, Linear Discriminant Analysis, K-Nearest Neighbors, Classification and Regression Trees, Naive Bayes, Support Vector Machine, XGBoost, Random Forest (RF), Gradient Boosting, AdaBoost, and Extra Tree, across six different data scaling methods: Normalization, Z-score Normalization, Min-Max Normalization, Max Normalization, Robust Scaler, and Quantile Transformer. However, they focused on only one dataset, the UCI - Heart Disease [52]. Despite that, their results are interesting; models based on Decision Trees showed the best performance without any scaling method, while K-Nearest Neighbors and Support Vector Machine achieved the lowest performance.
In a work on diabetes diagnosis using models such as Random Forest, Naive Bayes, K-Nearest Neighbors, Logistic Regression and Support Vector Machine [53] the researchers compared only three preprocessing scenarios: Normalization, Z-score Normalization, and no feature scaling. Their findings suggested that Random Forest, Naive Bayes, and Logistic Regression showed little sensitivity to these specific scaling approaches, and in some cases, their performance even worsened post-scaling.
Another problem is data leakage and the reproducibility of Machine Learning models. As shown in [36], many studies used preprocessing steps, such as feature scaling, on the entire dataset before splitting the data into training and test. Some studies also applied a scaling technique without knowing if that specific Machine Learning algorithm would benefit from it.
In [54] the authors demonstrate how feature scaling methods can impact the final model performance. Rather than relying on traditional scaling techniques, they propose a Generalized Logistic algorithm. This method showed particularly strong performance on datasets with a small number of samples, consistently outperforming models that used features scaled with Min-Max Normalization or Z-score Normalization.
The convergence of stochastic gradient descent is highly sensitive to the scale of input features, with studies such as [55] demonstrating that normalization is an effective method for improving convergence.
For leaf classification, a Neural Network called the Probabilistic Neural Network was used as a classifier and Min-Max normalization was applied [56]. For approval, Min-Max normalization was again used with a K-Nearest Neighbors classifier [57].
The impact of scaling and normalization techniques on NMR spectroscopic metabonomic datasets was examined in [58]. In a related study, [59] explored eight different normalization methods to enhance the biological interpretation of metabolomics data.
An in-depth study on the impact of data normalization on classification performance was presented in [41]. The authors evaluated fourteen normalization methods but employed only the K-Nearest Neighbor Classifier. Their findings indicate that normalization as a data preprocessing technique is affected by various data characteristics, such as features with differing statistical properties, the dominance of certain features, and the presence of outliers.
In an unsupervised task, [44] investigates the impact of the scaling of features on K-Means, highlighting its importance for datasets with features measured in different units. The study compares five scaling methods: Z-score Normalization, Min-Max Normalization, Percentile transformation, Maximum absolute scaling, and Robust Scaler. For cluster analysis, normalization helps prevent unwanted biases in external validation indices, such as the Jaccard, Fowlkes-Mallows, and Adjusted Rand Index, that may arise due to variations in the number of clusters or imbalances in class size distributions [60].
@@
In the context of glaucoma detection based on a combination of texture and higher-order spectral features [61], the authors demonstrate that Z-score normalization, when paired with a Random Forest classifier, achieves superior performance compared to a Support Vector Machine.
A most recent study [62] evaluates five Machine Learning models, but focuses exclusively on classification problems and applies only two feature scaling techniques. However, most of these studies do not explain the rationale for choosing these techniques. In addition, some apply normalization to the entire dataset before splitting it into training and testing sets, leading to data leakage.
III Methodology
@@
III-A Dataset
Tables I and II provide detailed information about the datasets used in this study, including the number of features, instances, and classes. All features are numeric, represented as either int64 or float64 types. The classification tasks are either binary or multi-class.
TABLE I: Datasets used for classification: Breast Cancer Wisconsin (Diagnostic) [63], Dry Bean Dataset [64], Glass Identification [65], Heart Disease [52], Iris [66], Letter Recognition [67], MAGIC Gamma Telescope [68], Rice (Cammeo and Osmancik) [69], and Wine [70].
@@ III-A Dataset
TABLE II: Datasets used for regression: Air Quality [71], Abalone [72], Appliances Energy Prediction [73], Concrete Compressive Strength [74], Forest Fires [75], Real Estate Valuation [76], and Wine Quality [77].
@@
III-B Train and Test set
To preserve the integrity of our analysis and avoid data leakage, we split the dataset into training and test sets prior to applying any preprocessing steps, such as feature scaling. Following standard practice in machine learning [78, 79], 70% of the data is allocated to the training set and 30% to the test set. While determining the optimal train-test split is inherently challenging in machine learning [80], our choice reflects a balance between reproducibility and the practical constraints posed by the relatively small size of some datasets. This ratio allows for both effective model training and reliable performance evaluation.
@@ III-B Train and Test set
III-C Feature Scaling Techniques
III-C1 Min-Max Normalization (MM)
$X_{\text{norm}}=\frac{X-X_{\text{min}}}{X_{\text{max}}-X_{\text{min}}},$ (1)
III-C2 Max Normalization (MA)
$X_{\text{norm}}=\frac{X}{max(~{}|X|~{})}$ (2)
III-C3 Z-score Normalization (ZSN)
$X_{\text{norm}}=\frac{X-\mu}{\sigma},$ (3)
III-C4 Variable Stability Scaling (VAST)
Variable stability scaling adjusts the data based on the stability of each feature. It is particularly useful for high-dimensional datasets and can be seen as a variation of standardization that incorporates the Coefficient of Variation (CV), $\frac{\mu}{\sigma}$, as a scaling factor [59]:
$X_{\text{norm}}=\frac{(X-\mu)}{\sigma}\frac{\mu}{\sigma}$ (4)
III-C5 Pareto Scaling (PS)
$X_{\text{norm}}=\frac{X-\mu}{\sqrt{\sigma}}$ (5)
III-C6 Mean Centered (MC)
$X_{\text{norm}}=X-\mu$ (6)
III-C7 Robust Scaler (RS)
$X_{\text{norm}}=\frac{X-X_{\text{median}}}{\text{IQR}}$
This method is robust to outliers [88].
III-C8 Quantile Transformation (QT)
Quantile transformation maps the data to a uniform or normal distribution. It is useful for non-linear data [88].
III-C9 Decimal Scaling Normalization (DS)
$X_{\text{norm}}=\frac{X}{10^{j}}$ (7)
where $j$ is the smallest integer such that $max(~{}|X_{norm}|~{})<1$
III-C10 Tanh Transformation (TT)
III-C11 Logistic Sigmoid Transformation (LS)
III-C12 Hyperbolic Tangent Transformation (HT)
III-D Machine Learning Algorithms
The following models were used:
  • Support Vector Machine (SVM) & Support Vector Regressor (SVR): With a linear kernel, this model finds the hyperplane that maximizes the margin between classes and for regression finds a function within a tolerance margin using the radial basis function (RBF) kernel [95, 96, 97].
  • K-Nearest Neighbors (KNN): An instance-based learning algorithm that classifies samples based on the most frequent label among their nearest neighbors, also have its regression variant that predicts the target by averaging the outputs of the $k$ nearest neighbors in the feature space [112, 113, 114].
III-E Metrics
III-E1 Classification Metrics
    $\text{Accuracy}=\frac{TP+TN}{TP+TN+FP+FN},$ (11)
    where $TP$, $TN$, $FP$, and $FN$ represent true positives, true negatives, false positives, and false negatives, respectively.
III-E2 Regression Metrics
    $\text{MAE}=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|$ (12)
    $\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$ (13)
III-E3 Computational Metrics
  • Memory Usage: measure memory usage during the scaling step.
  • Training Time: the time taken to train each model on a given dataset.
III-F Experiment Workflow
@@
III-G Python Script Descriptions
@@
Figure 1: Experimental design
III-H Source Code of this Experiments
For complete transparency and reproducibility, the source code, all experimental results, and detailed model parameters are publicly available in GitHub

### v4

@@ Abstract
This study addresses the lack of comprehensive evaluations of feature scaling by systematically assessing 12 techniques, including less common methods such as VAST and Pareto, in 14 machine learning algorithms and 16 datasets covering both classification and regression tasks. The impact of feature scaling was evaluated in terms of predictive performance (accuracy, MAE, MSE, R²) and computational efficiency (training time, inference time, memory usage). The results show that the ensemble methods (Random Forest, XGBoost, CatBoost, LightGBM) remain robust regardless of scaling, while models such as Logistic Regression, SVM, TabNet, and MLP are highly sensitive to the chosen scaler. By making all codes, results, and model parameters publicly available, this work provides reproducible, model-specific guidance for selecting scaling strategies in practical machine learning applications.
@@ I Introduction
Machine learning progress has been remarkable in several domains of knowledge engineering, notably driven by the rise of big data [1, 2], its applications in healthcare [3, 4, 5, 6], forecasting [7, 8, 9, 10], precision agriculture [11], wireless sensor networks [12], language tasks [13, 14, 15] and many other domains [16, 17, 18, 19, 20, 21]. All different applications compose the field of Machine Learning [14], which has become a major subarea of computer science and statistics due to its crucial role in the modern world [22]. Although these methods hold immense potential for the advancement of predictive modeling, their improper application has introduced significant obstacles [23, 24, 25].
@@
Consequently, applying a scaling method without a careful evaluation of its suitability for the specific problem and model may not be advisable and could negatively impact results. This practice risks compromising the validity of claims regarding model performance and may increase the risk of overfitting, that is, when a model adapts too closely to noise or idiosyncrasies in the training data, thereby reducing its generalization capability[30].
@@
The primary objective of this study is to evaluate the impact of different data scaling methods on the training process and performance metrics of various Machine Learning algorithms across multiple datasets. We employ 14 widely used Machine Learning models for tabular data, including Linear Regression, Logistic Regression, Support Vector Machines (SVMs), K-Nearest Neighbors (KNN), Multilayer Perceptron, Random Forest, TabNet, Naive Bayes, Classification and Regression Trees (CART), Gradient Boosting Trees, AdaBoost, LightGBM, CatBoost, and XGBoost. These models were evaluated using 12 different data scaling techniques, in addition to a baseline without scaling, across 16 datasets covering both classification and regression tasks. The selected models represent the state of the art in tabular data analysis, offering a favorable balance between predictive performance and computational efficiency, often outperforming deep learning techniques in this context [37, 38, 39, 40].
This study contributes by being the first to address the lack of comprehensive evaluation in feature scaling and algorithm benchmarking:
  • Testing 12 feature scaling techniques, including less common methods such as VAST, Pareto, Logistic Sigmoid, and Hyperbolic Tangent.
  • Benchmarking across 14 supervised learning algorithms, covering both regression and classification tasks.
  • Employing 16 diverse datasets from different domains to ensure generalizability of the findings.
  • Providing open-source code, experimental data, and parameters to ensure full reproducibility.
In Section II we cover some related work in similar studies. In Section III we explain more about each algorithm, each feature scaling technique and valuation metrics while Section IV, the study diagram and how the models were trained. In Section V we represent the final results of this study and some discussion. The limitations of our study are discussed in Section VI. Lastly, in Section VII we give a final conclusion of the current experiments and future works.
@@ II Related Work
For some Machine Learning models, feature scaling is extremely necessary, such as K-Nearest Neighbors[41], Neural Networks[42, 43, 44], and SVMs[45, 46].
Despite its fundamental role in Machine Learning pipelines, the impact of feature scaling remains an underexplored area in the literature. Most existing studies examine only a limited number of algorithms or datasets, and often provide minimal analysis of the specific effects of different scaling techniques[47] on each particular algorithm’s performance. Studies that comprehensively evaluate various scaling methods across a broad range of models and datasets, such as the approach taken in this work, are scarce. In many Machine Learning cases, preprocessing is briefly mentioned, with scaling treated as a routine step rather than a variable worthy of in-depth investigation.
@@
In[48], the authors compared 6 normalization methods applied only in an SVM classifier used to improve intrusion data. In the context of glaucoma detection based on a combination of texture and higher-order spectral features[49], the authors demonstrate that Z-score normalization, when paired with a Random Forest classifier, achieves superior performance compared to a Support Vector Machine.
In[50], the authors demonstrate how feature scaling methods can impact the final model performance; however, only 2 algorithms were used in a binary classification task. A more complete work[51] evaluates 12 learning algorithms and 6 different normalization techniques.
An in-depth study on the impact of data normalization on classification performance was presented in[52], where the authors evaluated 14 normalization methods but employed only the K-Nearest Neighbor Classifier. Another study also used K-Nearest Neighbor Classifier and two SVMs[53] with 7 scaling techniques, but only in 1 dataset.
In[54], the authors evaluated 11 Machine Learning algorithms, across 6 different data scaling methods. However, they focused on only 1 dataset, the UCI - Heart Disease[55].
In a study on a single dataset (diabetes diagnosis), researchers tested models including Random Forest, Naive Bayes, KNN, Logistic Regression, and SVM[56], but compared only 3 preprocessing scenarios.
A recent study[57] evaluated 5 Machine Learning models, focusing solely on classification. However, some approaches normalized the entire dataset before splitting, introducing data leakage.
TABLE I: Comparative summary of related work in feature scaling.
Paper Year Scalers Algorithms Datasets Gap
[48] 2011 6 1 1
Limited to
1 algorithm
and 1 dataset
[49] 2011 2 4 1
Limited to
2 scalers
and 1 dataset
[50] 2016 3 2 16
Only 2 algo-
rithms and binary
classification
[51] 2017 2 12 1
Limited to
2 scalers
and 1 dataset
[52] 2019 14 1 21
Focused only
on 1 algorithm
[53] 2020 6 3 1
Limited to
3 algorithms and
only 1 dataset
[54] 2021 6 11 1
Only 1 dataset
[56] 2022 3 5 1
Restricted to
1 dataset
and 3 scalers
[57] 2024 2 5 1
Only 2 algo-
rithms, 1 dataset,
and scaler applied
before splitting
Compared to prior studies (see Table I) our work offers a broader and more systematic evaluation of feature scaling. We benchmark 12 techniques across 14 algorithms and 16 datasets. Unlike earlier research, we ensure reproducibility by releasing open-source code, parameters, and data.
III Background
III-A Feature Scaling Techniques
We selected 12 feature scaling techniques to balance breadth, diversity, and experimental feasibility across the 14 algorithms and 16 datasets considered. This set includes widely adopted methods such as Min–Max, Standardization, and Robust Scaling, which represent standard practice in Machine Learning, alongside less commonly benchmarked but theoretically relevant approaches such as VAST, Pareto, Logistic Sigmoid, and Hyperbolic Tangent, which offer distinctive transformations with potential impact on model performance. While additional scalers exist, including too many would make the experiments intractable and the results harder to interpret. By focusing on these 12 methods, we provide a comprehensive yet manageable benchmark that captures both established practices and underexplored alternatives, ensuring the study’s value as a practical reference for future work.
III-A1 Min-Max Normalization (MM)
$X_{\text{norm}}=\frac{X-X_{\text{min}}}{X_{\text{max}}-X_{\text{min}}},$ (1)
III-A2 Max Normalization (MA)
$X_{\text{norm}}=\frac{X}{max(~|X|~)}$ (2)
III-A3 Z-score Normalization (ZSN)
$X_{\text{norm}}=\frac{X-\mu}{\sigma},$ (3)
III-A4 Variable Stability Scaling (VAST)
Variable stability scaling adjusts the data based on the stability of each feature. It is particularly useful for high-dimensional datasets and can be seen as a variation of standardization that incorporates the Coefficient of Variation (CV), $\frac{1}{CV}=\frac{\mu}{\sigma}$, as a scaling factor [61, 62]:
$X_{\text{norm}}=\frac{(X-\mu)}{\sigma}\frac{\mu}{\sigma}$ (4)
III-A5 Pareto Scaling (PS)
$X_{\text{norm}}=\frac{X-\mu}{\sqrt{\sigma}}$ (5)
III-A6 Mean Centered (MC)
$X_{\text{norm}}=X-\mu$ (6)
III-A7 Robust Scaler (RS)
$X_{\text{norm}}=\frac{X-X_{\text{median}}}{\text{IQR}}$
This method is robust to outliers [67, 68].
III-A8 Quantile Transformation (QT)
Quantile transformation maps the data to a uniform or normal distribution. It is useful for non-linear data [69, 70].
III-A9 Decimal Scaling Normalization (DS)
$X_{\text{norm}}=\frac{X}{10^{j}}$ (7)
where $j$ is the smallest integer such that $max(~|X_{norm}|~)<1$
III-A10 Tanh Transformation (TT)
III-A11 Logistic Sigmoid Transformation (LS)
III-A12 Hyperbolic Tangent Transformation (HT)
III-B Machine Learning Algorithms
The following models were used:
  • Support Vector Machine (SVM) & Support Vector Regressor (SVR): This model finds the hyperplane that maximizes the margin between classes and for regression finds a function within a tolerance margin using the radial basis function (RBF) kernel[77, 78, 79].
  • K-Nearest Neighbors (KNN): An instance-based learning algorithm that classifies samples based on the most frequent label among their nearest neighbors, also has a regression variant that predicts the target by averaging the outputs of the $k$ nearest neighbors in the feature space [94, 95, 96].
III-C Metrics
III-C1 Classification Metrics
    $\text{Accuracy}=\frac{TP+TN}{TP+TN+FP+FN},$ (11)
    where $TP$, TNTN, FPFP, and $FN$ represent true positives, true negatives, false positives, and false negatives, respectively.
III-C2 Regression Metrics
    $\text{MAE}=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|$ (12)
    $\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$ (13)
III-C3 Computational Metrics
  • Memory Usage: Measure the increase in memory usage during dataset loading and the feature scaling step.
  • Training Time: The time taken to train each model on a given dataset.
IV Methodology
@@
IV-A Dataset
Tables II and III provide detailed information about the datasets used in this study, including the number of features, instances, and classes. All features are numeric, represented as either int64 or float64 types. The classification tasks are either binary or multi-class.
TABLE II: Datasets used for classification: Breast Cancer Wisconsin (Diagnostic) [101], Dry Bean Dataset [102], Glass Identification [103], Heart Disease [55], Iris [104], Letter Recognition [105], MAGIC Gamma Telescope [106], Rice (Cammeo and Osmancik) [107], and Wine [108].
@@ IV-A Dataset
TABLE III: Datasets used for regression: Air Quality [109], Abalone [110], Appliances Energy Prediction [111], Concrete Compressive Strength [112], Forest Fires [113], Real Estate Valuation [114], and Wine Quality [115].
@@
IV-B Train and Test set
To preserve the integrity of our analysis and avoid data leakage, we split the dataset into training and test sets prior to applying any preprocessing steps, such as feature scaling.
@@ IV-B Train and Test set
Following standard practice in machine learning[116],[117], 70% of the data was allocated to the training set and 30% to the test set. There is no universal optimal train–test split in Machine Learning[118]. The choice depends on dataset size, the bias–variance trade-off, and computational constraints. In this study, we adopted a 70/30 split as a practical balance that allows sufficient data for model training while still allowing reliable performance evaluation. To ensure reproducibility, random seeds were fixed in all experiments.
IV-C Experiment Workflow
@@
IV-D Python Script Descriptions
@@
Figure 1: Experimental design.
IV-E Source Code of this Experiments
To ensure full transparency and reproducibility, we provide the source code, complete experimental results, statistical analyses, detailed model parameters, and all accompanying plots and figures. They are publicly available in GitHub.